Problem 5 15 points Your lawn care service is considering two commercial mowers for purchase, as indicated below. Assume thatAnswer is in bold. Need to know how to work the problem. Thanks!
Answer #1

Repeat small mowers cash flow twice

PW (small) = -2000 - 2000(P/F, 12%, 5) + 1000(P/A, 12%, 10) + 500(P/F, 12%, 5) + 500(P/F, 12%, 10)

= -2000 - 2000*1.12^-5 + 1000*5.6502 + 500*1.12^-5 + 500*1.12^-10

= 2960.05

PW (large) = -5000 + 1500(P/A, 12%, 10) + 700(P/F, 12%, 10)

= -5000 + 1500*5.6502 + 700*1.12^-10

= 3700.68

Now annual worth is found as

AW(small) = PW(small)/(P/A, 12%, 10)

= 2960.05/5.6502

= 523.89

AW(large) = PW(large)/(P/A, 12%, 10)

= 3700.68/5.6502

= 654.97

Select large mower.